OpenStudy (anonymous):

Find the number halfway between 3/5 and 7/10 . 27/40 2/3 7/10 13/20

OpenStudy (alexandervonhumboldt2):

what you think>?

OpenStudy (tkhunny):

Have you considered the average? \(\dfrac{\dfrac{3}{5}+\dfrac{7}{10}}{2}\) Have you considered getting a matching denominator: \(\dfrac{3}{5}\;vs\;\dfrac{7}{10}\;suggests\;\dfrac{6}{10}\;vs\;\dfrac{7}{10}\;or\;\dfrac{12}{20}\;vs\;\dfrac{14}{10}\) Now, WHY would I go to the trouble to do the 20ths?

OpenStudy (mathstudent55):

Of course, you meant to write \(\dfrac{14}{20} \), not \(\dfrac{14}{10} \).

OpenStudy (tkhunny):

Right. I fixed that twice. I guess it took three.